Use this chart for the first three questions. Consider these points from a PPF:

Wheat              210      150      145      135      0

Corn                0          20        21        22        150

The opportunity cost of 210 units of wheat is:

The opportunity cost of 21 units of corn is equal to:

The marginal opportunity cost of the 22nd unit of corn is equal to:
Answer #1

Opportunity cost is the best benefit foregone if an alternative is not chosen.

210 units:

Answer: 150 units of C

Opportunity cost of W is the units of corn not produced, which should be 150.

21 units:

Answer: 65 units of W

This is the difference of units sacrificed of W for an increasing production of C from 0 units to 21 units.

Opportunity cost = 145 – 210 = 65 units (ignoring sign)

22 units:

Answer: 10 units of W

This is the difference sacrificed of W that has to be divided by the difference in C.

Marginal opportunity cost = (135 – 145) / (22 – 21) = - 10 / 1 = 10 units (ignoring sign).